# Investigating the heterogenous effects of a massive content moderation intervention via Difference-in-Differences

### v3

@@ 3 Dataset
Baseline group. We obtain accurate causal estimates of the effects of The Great Ban by comparing the behavioral changes exhibited by the core users to possible changes exhibited during the same time by a set of reference Reddit users who were not affected by the ban. To reach this goal, we first identify a small set of related, but non-banned, subreddits to act as a baseline for the banned ones. To obtain this baseline, we first preprocessed every comment in IN-BEFORE for each banned subreddit by removing URLs, stopwords, and words less than four characters long. Subsequently, we performed stemming. For r/chapotraphouse we executed these steps only on a random sample of 1M comments, due to the sheer number of comments in the subreddit. The preprocessed comments were then fed to BERTopic  (Grootendorst, 2022) to extract up to ten relevant topics, each consisting of ten keywords. We then used the 100 descriptive keywords for each subreddit to project it with BERT (Devlin et al., 2019) into a 768-dimensional embedding space. This representation is suitable to highlight commonalities between the banned subreddits, such as via clustering, which we leverage to identify the small set of baseline subreddits. However, the high dimensionality of the embedding space is not suitable to directly perform clustering (Domingos, 2012; Zimek, 2018). Given the low number of data points – the 15 banned subreddits – we therefore first greatly reduce the dimensionality with PCA before clustering the banned subreddits with k-means (Ikotun et al., 2023). We experimented with multiple numbers $p$ of principal components and values for $k$, ultimately choosing $p=5$ and $k=5$ by optimizing the silhouette coefficient. Table 2 illustrates the clustering output. We now identify a set of reference non-banned subreddits for each cluster of banned subreddits, with the exception of cluster 5 that is formed by particularly small subreddits (i.e., r/imgoingtohellforthis2 and r/wojak). For clusters one to four, we computed the cluster centroid and found the most representative keywords as those having the smallest Euclidean distance to the centroid. We then retrieved all comments posted during the pre-ban period that contained those keywords and identified the non-banned subreddits where such keywords occurred more frequently. These subreddits are related to the non-banned ones and are therefore suitable candidates for representing a baseline for our causal analyses. As reported in Table 2, r/politics, r/askreddit, and r/worldnews are overall the three most representative non-banned subreddits for our clusters. As reported in Table 1, given the numerosity of the baseline subreddits, we selected a random stratified sample of 17K users, which is comparable to the number of core users in the focus group of banned subreddits. From this set, we removed 45 users who posted at least one comment in one of the 15 banned subreddits, so as to avoid any overlap between the focus and baseline groups.
[…]
270 aband. before (ABA) remain. before (BEF) remain. after (AFT) DiD subreddit users mean stdev users mean stdev mean stdev ABA vs BEF BEF vs AFT r/chapotraphouse 886 0.163 0.123 8,319 0.145 0.084 0.138 0.094 $-0.018$** $-0.005$** r/the_donald 987 0.163 0.122 3,145 0.145 0.081 0.145 0.113 $-0.018$** $-0.005$** r/darkhumorandmemes 225 0.183 0.108 1,392 0.161 0.073 0.154 0.098 $-0.022$** $-0.009$** r/consumeproduct 444 0.190 0.090 1,275 0.158 0.045 0.160 0.099 $-0.032$** $-0.009$** r/gendercritical 333 0.184 0.119 706 0.167 0.071 0.141 0.116 $-0.017$** $-0.019$** r/thenewright 151 0.177 0.092 575 0.153 0.076 0.145 0.098 $-0.024$** $-0.009$** r/soyboys 162 0.200 0.074 432 0.177 0.072 0.163 0.087 $-0.023$** $-0.011$** r/shitneoconssay 171 0.190 0.084 384 0.165 0.071 0.160 0.098 $-0.025$** $-0.007$** r/debatealtright 148 0.171 0.084 328 0.153 0.070 0.155 0.115 $-0.018$* $-0.002$* r/darkjokecentral 38 0.221 0.154 270 0.155 0.070 0.144 0.092 $-0.066$** $-0.005$** r/wojak 70 0.192 0.068 170 0.160 0.065 0.143 0.044 $-0.032$** $-0.011$** r/hatecrimehoaxes 45 0.175 0.073 143 0.166 0.074 0.158 0.111 $-0.009$ $-0.004$** r/ccj2 26 0.188 0.093 119 0.133 0.076 0.125 0.106 $-0.055$** $-0.008$** r/imgoingtohellforthis2 20 0.202 0.144 72 0.169 0.077 0.187 0.141 $-0.033$ $-0.004$* r/oandaexclusiveforum 4 0.129 0.107 55 0.188 0.086 0.182 0.092 $+0.059$ $-0.010$** focus group overall 2,577 0.167 0.120 13,963 0.147 0.083 0.141 0.101 $-0.020$** $-0.006$** r/askreddit 274 0.146 0.093 15,576 0.121 0.066 0.120 0.072 $-0.025$** $-0.001$** r/worldnews 143 0.146 0.083 9,907 0.125 0.067 0.124 0.072 $-0.021$** $-0.001$** r/politics 136 0.146 0.079 8,862 0.125 0.068 0.124 0.072 $-0.021$** $-0.001$** baseline group overall 304 0.145 0.096 16,633 0.120 0.067 0.119 0.073 $-0.025$** $-0.001$** *: $p<0.1$; **: $p<0.05$; ***: $p<0.01$ Table 3: Subreddit-wise mean toxicity scores for users who left Reddit after the ban (ABA) and for the remaining ones. For the latter, toxicity scores were computed both before (BEF) and after (AFT) the ban. The ABA vs BEF and BEF vs AFT columns show the effect sizes and the statistical significance of the differences in toxicity. The causal effect of The Great Ban on the remaining users was computed via Difference-in-Differences (DiD), by discounting the effects measured for the banned subreddits with those of the baseline (non-banned) subreddits.
@@ 4.1 RQ1: Effectiveness of The Great Ban
Abandoning users. Table 1 underlines a difference of 2,577 users between the OUT-BEFORE and OUT-AFTER datasets, meaning that (15.6%) of the users became inactive following The Great Ban. The percentage of users who have abandoned the baseline subreddits is much smaller, at approximately 2%, corresponding to 304 users. Since there is no sign of activity from these users during the seven months following the ban, we can presume that they not only abandoned Reddit but also likely migrated to other platforms. This is the first clear effect of The Great Ban. To provide more details, for each subreddit we compare the toxicity of abandoning users to the toxicity of those who remained on the platform. As shown in Table 3, in 14 out of 15 subreddits, the abandoning users exhibited higher pre-ban toxicity than the others. This means that toxic users were more inclined to abandon the platform after the ban, compared to their less toxic fellows. In 12 subreddits, the difference in toxicity between users who abandoned and those who did not is statistically significant ($p<0.05$) according to a t-test for unpaired data. Additionally, abandoning users exhibited higher standard deviation, suggesting greater variability in toxicity.
@@
Remaining users. Besides leading a subset of users to abandon the platform, the ban may have also contributed to shifts in toxicity among the remaining users. Table 3 presents, for each subreddit, toxicity scores for the set of remaining users, both before and after the ban. It shows how, for the most part, the remaining users reduced their toxicity scores after the ban. In particular, the scores aggregated by subreddit reveal that users from all 15 subreddits exhibited a slight decrease in their toxicity. We performed a Difference-in-Differences (DiD) analysis to compare the changes with the baseline group, and proved how this decrease is also statistically significant ($p<0.05$). The effect sizes obtained comparing remaining users before and after the ban are notably smaller than those obtained comparing the abandoning users to remaining ones. In fact, after The Great Ban, there was an average reduction in overall toxicity of 4.08% — a relatively modest decrease. Furthermore, in 14 of the 15 subreddits, the standard deviation for toxicity values is higher after the ban compared to before. This indicates that the ban has increased variability in the behavior of the remaining users, a phenomenon that we explore more thoroughly with further analyses.
@@
User-level effects. Until now, we presented results at community-level, showing that after the ban, 15.6% of core users abandoned the platform and the remaining ones reduced slightly their average toxicity by 4.08%. To provide a more in-depth analysis, we also examine user-level effects by analyzing the changes in toxicity experienced by each of the remaining 13,963 users. The central panel of Figure 2 shows a slope chart representing user-level toxicity changes across all users, regardless of the subreddit in which they were active. A single line is associated with each user and line slopes represent the amount of increase or decrease in toxicity. Rising lines can be colored in different shades of red based on their slope, representing users who became more toxic after the ban. Instead, the decreasing lines are in different shades of blue and represent users who reduced their toxicity. Figure 2 also features marginal boxplots displaying the toxicity distributions of remaining users before (left side of the slope chart) and after (right side) the ban. In particular, the leftmost boxplot displays the toxicity distribution for users who left the platform. Finally, the panel at the bottom shows the distribution of the changes in user-level toxicity in the form of a beeswarm plot. This is useful to identify outliers and study both tails of the distribution – those associated with significant toxicity increases (red dots on the right-hand side of the beeswarm plot) and decreases (blue dots on the left-hand side).
@@
The three boxplots in Figure 2 align with and support the general toxicity trends detailed in Table 3. Users who abandoned the platforms have the largest average toxicity with respect to those who stayed. The group displaying the lowest average toxicity is the one formed by the remaining users after the ban. The slope chart in Figure 2 shows that most lines have relatively small positive or negative slopes. This means that most users manifested little changes in toxicity, whether it is a change in positive or negative. At the same time, the slope chart also displays a notable number of steep lines, representing users who showed significant changes in toxicity. Specifically, the steep red lines outnumber the blue ones. This means that the majority of users who drastically changed their toxicity became more toxic. This finding qualitatively illustrates an unintended consequence of The Great Ban: it led a significant minority of users to become resentful, resulting in markedly more toxic behavior. The points where the lines intersect the y axis on the right side of the plot indicate the users’ toxicity levels after the ban. This is also shown in the boxplot on the right side of the slope chart. As illustrated in Figure 2 and as suggested by the standard deviation values in Table 3, there is increased variability in user toxicity following the ban. What led to this increase in variability is the presence of the aforementioned resentful users.
@@
Figure 2 presents the aggregated results across all subreddits. Nonetheless, the same plots can be used to investigate the effects of The Great Ban among single subreddits, as done by Cima et al. (2024b). This is useful for identifying common patterns and potential differences among the subreddits. We therefore conducted the analysis separately for users in each subreddit. The comparison across different subreddits reaffirmed earlier findings that most users who experienced significant toxicity changes after the ban increased their toxicity. However, this behavior was more pronounced in some subreddits and less evident in others. Overall, we observed that the findings for RQ1 in  (Cima et al., 2024b), which utilized the median, are similar to those presented in this section when using the mean and a causal inference method like DiD.

### v4

@@ 3.2 Baseline group
We obtain accurate causal estimates of the effects of The Great Ban by comparing the behavioral changes exhibited by the core users to possible changes exhibited during the same time by a set of reference Reddit users who were not affected by the ban. To reach this goal, we first identify a small set of related, but non-banned, subreddits to act as a baseline for the banned ones. To obtain this baseline, we first preprocessed every comment in IN-BEFORE for each banned subreddit by removing URLs, stopwords, and words less than four characters long. Subsequently, we performed stemming. For r/chapotraphouse we executed these steps only on a random sample of 1M comments, due to the sheer number of comments in the subreddit. The preprocessed comments were then fed to BERTopic  [26] to extract up to ten relevant topics, each consisting of ten keywords. We then used the 100 descriptive keywords for each subreddit to project it with BERT [18] into a 768-dimensional embedding space. This representation is suitable to highlight commonalities between the banned subreddits, such as via clustering, which we leverage to identify the small set of baseline subreddits. However, the high dimensionality of the embedding space is not suitable to directly perform clustering [20, 64]. Given the low number of data points – the 15 banned subreddits – we therefore first greatly reduce the dimensionality with PCA before clustering the banned subreddits with k-means [33]. We experimented with multiple numbers $p$ of principal components and values for $k$, ultimately choosing $p=5$ and $k=5$ by optimizing the silhouette coefficient. Table 2 illustrates the clustering output. We now identify a set of reference non-banned subreddits for each cluster of banned subreddits, with the exception of cluster 5 that is formed by particularly small subreddits (i.e., r/imgoingtohellforthis2 and r/wojak). For clusters one to four, we computed the cluster centroid and found the most representative keywords as those having the smallest Euclidean distance to the centroid. We then retrieved all comments posted during the pre-ban period that contained those keywords and identified the non-banned subreddits where such keywords occurred more frequently. These subreddits are related to the non-banned ones and are therefore suitable candidates for representing a baseline for our causal analyses. As reported in Table 2, r/politics, r/askreddit, and r/worldnews are overall the three most representative non-banned subreddits for our clusters. As reported in Table 1, given the numerosity of the baseline subreddits, we selected a random stratified sample of 17K users, which is comparable to the number of core users in the focus group of banned subreddits. From this set, we removed 45 users who posted at least one comment in one of the 15 banned subreddits, so as to avoid any overlap between the focus and baseline groups.
Selecting an appropriate baseline group presents inherent challenges, as different approaches come with trade-offs. While our method favors larger subreddits due to keyword-based selection, this choice ensures that the reference subreddits provide a stable and representative sample of general trends on the platform. An alternative approach, such as selecting smaller but thematically closer subreddits, could have introduced significant issues. For instance, r/donaldtrump – a seemingly relevant alternative for r/The_Donald, had minimal activity before the Great Ban and only surged afterward, making it both temporally inconsistent and dependent on the studied event. As a result, it would not serve as a valid counterfactual. Our selected baseline subreddits, while only approximating an ideal control group, strike a balance between ensuring thematic relevance and maintaining independence from the intervention, thus allowing for meaningful causal inference.
3.3 Toxicity
Several automated methods exist for assessing the toxicity of textual content. In this study, we compute toxicity scores using Detoxify [28], an open-source deep learning classifier that provides reliable toxicity assessments. Detoxify achieves performance comparable to state-of-the-art solutions [13] and can be installed and executed locally. Due to its effectiveness and accessibility, it has been widely adopted in recent research on online toxicity [22, 41]. Among the various scores provided by Detoxify, we specifically use the toxicity score, as it has been shown to exhibit the highest correlation with toxicity assessments from other established methods [13].
aband. before (ABA) remain. before (BEF) remain. after (AFT)
subreddit users mean stdev users mean stdev mean stdev
r/chapotraphouse 886 0.163 0.123 8,319 0.145 0.084 0.138 0.094
r/the_donald 987 0.163 0.122 3,145 0.145 0.081 0.145 0.113
r/darkhumorandmemes 225 0.183 0.108 1,392 0.161 0.073 0.154 0.098
r/consumeproduct 444 0.190 0.090 1,275 0.158 0.045 0.160 0.099
r/gendercritical 333 0.184 0.119 706 0.167 0.071 0.141 0.116
r/thenewright 151 0.177 0.092 575 0.153 0.076 0.145 0.098
r/soyboys 162 0.200 0.074 432 0.177 0.072 0.163 0.087
r/shitneoconssay 171 0.190 0.084 384 0.165 0.071 0.160 0.098
r/debatealtright 148 0.171 0.084 328 0.153 0.070 0.155 0.115
r/darkjokecentral 38 0.221 0.154 270 0.155 0.070 0.144 0.092
r/wojak 70 0.192 0.068 170 0.160 0.065 0.143 0.044
r/hatecrimehoaxes 45 0.175 0.073 143 0.166 0.074 0.158 0.111
r/ccj2 26 0.188 0.093 119 0.133 0.076 0.125 0.106
r/imgoingtohellforthis2 20 0.202 0.144 72 0.169 0.077 0.187 0.141
r/oandaexclusiveforum 4 0.129 0.107 55 0.188 0.086 0.182 0.092
focus group overall 2,577 0.167 0.120 13,963 0.147 0.083 0.141 0.101
r/askreddit 274 0.146 0.093 15,576 0.121 0.066 0.120 0.072
r/worldnews 143 0.146 0.083 9,907 0.125 0.067 0.124 0.072
r/politics 136 0.146 0.079 8,862 0.125 0.068 0.124 0.072
baseline group overall 304 0.145 0.096 16,633 0.120 0.067 0.119 0.073
*: $p<0.1$; **: $p<0.05$; ***: $p<0.01$
Table 3: Subreddit-wise mean and standard deviation of toxicity scores for users who left Reddit after the ban (ABA) and for those who remained. For the latter, toxicity scores were computed both before (BEF) and after (AFT) the ban.
toxicity mean toxicity stdev
subreddit ABA vs BEF BEF vs AFT ABA vs BEF BEF vs AFT
r/chapotraphouse $-0.018$** $-0.005$** $-0.047$** $+0.012$**
r/the_donald $-0.018$** $-0.005$** $-0.046$** $+0.040$**
r/darkhumorandmemes $-0.022$** $-0.009$** $-0.047$** $+0.030$**
r/consumeproduct $-0.032$** $-0.009$** $-0.024$** $+0.029$**
r/gendercritical $-0.017$** $-0.019$** $-0.029$** $+0.023$**
r/thenewright $-0.024$** $-0.009$** $-0.018$** $+0.026$**
r/soyboys $-0.023$** $-0.011$** $-0.020$ $+0.012$**
r/shitneoconssay $-0.025$** $-0.007$** $-0.010$ $+0.031$**
r/debatealtright $-0.018$* $-0.002$* $-0.017$** $+0.052$**
r/darkjokecentral $-0.066$** $-0.005$** $-0.103$ $+0.024$**
r/wojak $-0.032$** $-0.011$** $-0.005$ $+0.044$**
r/hatecrimehoaxes $-0.009$ $-0.004$** $+0.020$ $+0.015$*
r/ccj2 $-0.055$** $-0.008$** $-0.018$ $+0.033$**
r/imgoingtohellforthis2 $-0.033$ $-0.004$* $-0.017$** $+0.065$**
r/oandaexclusiveforum $+0.059$ $-0.010$** $-0.024$ $+0.009$
focus group overall $-0.020$** $-0.006$** $-0.044$** $-0.077$**
r/askreddit $-0.025$** - $+0.033$** $+0.007$**
r/worldnews $-0.021$** - $+0.020$** $+0.006$**
r/politics $-0.021$** - $+0.016$** $+0.006$**
baseline group overall $-0.025$** - $+0.037$** $+0.007$**
* : $p<0.1$; ** : $p<0.05$; *** : $p<0.01$
Table 4: Effect sizes of The Great Ban and their statistical significance. Mean toxicity results for BEF vs AFT are obtained via Difference-in-Differences, while results for all other columns via direct subtraction. Statistical significance of the differences between mean values is obtained via t-tests, while that of differences in variability is obtained via F-tests on the variances.
@@ 4.1 RQ1: Effectiveness of The Great Ban
Abandoning users. Table 1 underlines a difference of 2,577 users between the OUT-BEFORE and OUT-AFTER datasets, meaning that (15.6%) of the users became inactive following The Great Ban. The percentage of users who have abandoned the baseline subreddits is much smaller, at approximately 2%, corresponding to 304 users. Since there is no sign of activity from these users during the seven months following the ban, we can presume that they not only abandoned Reddit but also likely migrated to other platforms. This is the first clear effect of The Great Ban. To provide more details, for each subreddit we compare the toxicity of abandoning users to the toxicity of those who remained on the platform. As shown in Table 3, in 14 out of 15 subreddits, the abandoning users exhibited higher pre-ban toxicity than the others. This means that toxic users were more inclined to abandon the platform after the ban, compared to their less toxic fellows. In 12 subreddits, the difference in toxicity between users who abandoned and those who did not is statistically significant ($p<0.05$) according to a t-test for unpaired data. Additionally, abandoning users exhibited significantly larger standard deviation, as reported in Table 4, suggesting greater variability in toxicity.
Checking model assumptions. To estimate the causal effect of The Great Ban, we rely on Difference-in-Differences (DiD), which compares outcome changes over time between a treated and a control group. Here, we refer to the treated and control groups as the focus and baseline groups, respectively. This approach relies on the assumption that, without treatment, both groups would follow a common trend, allowing us to isolate the treatment effect. The following OLS regression model is used to estimate the effects:
$Y_{it}=\beta_{0}+\beta_{1}T_{t}+\beta_{2}S_{i}+\beta_{3}(T_{t}\cdot S_{i})+\epsilon_{it}$ (1)
where $T_{t}$ is a time indicator ($1$ after the intervention, 00 before), $S_{i}$ identifies the treatment group ($1$ if treated, 00 otherwise), and the interaction term $\beta_{3}$ captures the estimated effect. In practice, the DiD estimator for the effect can be computed as:
$\delta=(y_{22}-y_{12})-(y_{21}-y_{11})$ (2)
where $y_{st}$ represents the outcome for group $s$ at time $t$. Additionally, $t=1$ refers to the pre-ban period while $t=2$ refers to the post-ban period, and $s=1$ indicates the baseline group while $s=2$ indicates the focus group. In addition to the standard OLS assumptions, the DiD model in Eq. (1) relies on the Stable Unit Treatment Value Assumption (SUTVA), ensuring that no spillover effects exist between treated and control groups, and on the parallel trends assumption, which states that, in the absence of intervention, both groups would have followed similar trends over time. We assume SUTVA to be valid based on our choice of baseline subreddits, which are completely non-overlapping it terms of users with the banned subreddits. Furthemore, we empirically verified the parallel toxicity trends between the focus and baseline groups by analyzing the monthly pre-intervention toxicity levels, which confirmed the parallel trends.
Remaining users. Besides leading a subset of users to abandon the platform, the ban may have also contributed to shifts in toxicity among the remaining users. Table 3 presents, for each subreddit, toxicity scores for the set of remaining users, both before and after the ban. It shows how, for the most part, the remaining users reduced their toxicity scores after the ban. In particular, the scores aggregated by subreddit reveal that users from all 15 subreddits exhibited a slight decrease in their toxicity. Table 4 presents the size and statistical significance of the effects of the The Great Ban in terms of mean toxicity and its standard deviation, for all banned subreddits. Statistical significance of the differences between the means is obtained via t-tests, while that of dispersion is computed via F-tests on the variances. The results in Table 4 indicate a consistent decrease in mean toxicity between abandoning and remaining users across all banned subreddits, with most differences being statistically significant ($p<0.01$). This suggests that users who remained on Reddit exhibited lower toxicity levels than those who left after the ban. Differences in standard deviation show mixed effects. When comparing abandoning and remaining users before the ban, we note that the latter showed less variability. This result is statistically significant ($p<0.01$) in the majority of subreddits. Instead, when comparing remaining users before and after the ban, we find that the variability in toxicity increases after the intervention. This indicates that the ban has increased variability in the behavior of the remaining users, a phenomenon that we explore more thoroughly with subsequent analyses. We also note that the effect sizes measured when comparing remaining users before and after the ban are notably smaller than those obtained comparing the abandoning users to remaining ones before the ban. In fact, after The Great Ban, there was an average reduction in overall toxicity of 4.08% — a relatively modest decrease. To further validate our findings, we conducted the same analyses using the user-level frequency of toxic comments (i.e., those with toxicity scores $>0.5$) instead of the mean toxicity. We then compared these results with those reported in Tables 3 and 4, finding that the results remained largely consistent, which reinforces the observed effects.
@@
User-level effects. Until now, we presented results at community-level, showing that after the ban, 15.6% of core users abandoned the platform and the remaining ones reduced slightly their average toxicity by 4.08%. To provide a more in-depth analysis, we also examine user-level effects by analyzing the changes in toxicity experienced by each of the 13,963 remaining users. The central panel of Figure 2 shows a slope chart representing user-level toxicity changes across all users, regardless of the subreddit in which they were active. A single line is associated with each user and line slopes represent the amount of increase or decrease in average toxicity. Rising lines are colored in different shades of red based on their slope, representing users who became more toxic after the ban. Instead, the decreasing lines are in different shades of blue and represent users who reduced their toxicity. Figure 2 also features marginal boxplots displaying the toxicity distributions of remaining users before (left side of the slope chart) and after (right side) the ban. The leftmost boxplot displays the toxicity distribution for users who left the platform. Finally, the panel at the bottom shows the distribution of the changes in user-level toxicity in the form of a beeswarm plot. This is useful to identify outliers and study both tails of the distribution – those associated with significant toxicity increases (red dots on the right-hand side of the beeswarm plot) and decreases (blue dots on the left-hand side).
@@
The three boxplots in Figure 2 align with and support the general toxicity trends detailed in Table 3. Users who abandoned the platform have the largest average toxicity with respect to those who stayed. The group displaying the lowest average toxicity is the one formed by the remaining users after the ban. The slope chart in Figure 2 shows that most lines have relatively small positive or negative slopes. This means that most users manifested little changes in toxicity, whether it is a change in positive or negative. At the same time, the slope chart also displays a notable number of steep lines, representing users who showed significant changes in toxicity. Specifically, the steep red lines outnumber the blue ones. This means that the majority of users who drastically changed their toxicity became more toxic. This result is consistent across multiple toxicity indicators – including average toxicity, proportion of toxic comments, and absolute toxic comment volume – and remains valid even when focusing only on users with stable and substantial activity before and after the ban. This finding qualitatively illustrates an unintended consequence of The Great Ban: it led a significant minority of users to become resentful, resulting in markedly more toxic behavior. The points where the lines intersect the y axis on the right side of the plot indicate the users’ toxicity levels after the ban. This is also shown in the boxplot on the right side of the slope chart. As illustrated in Figure 2 and as suggested by the standard deviation values in Table 3, there is increased variability in user toxicity following the ban. What led to this increase in variability is the presence of the aforementioned resentful users. Additional correlation analyses further confirm that these users not only exhibited higher toxicity scores but also posted a higher raw number of toxic comments, reinforcing that the observed shifts reflect real behavioral changes rather than artifacts of reduced post-ban activity.
@@
Figure 2 presents the aggregated results across all subreddits. Nonetheless, the same plots can be used to investigate the effects of The Great Ban among single subreddits, as done in our earlier work [13]. This is useful for identifying common patterns and potential differences among the subreddits. We therefore conducted the analysis separately for users in each subreddit. The comparison across different subreddits reaffirmed earlier findings that most users who experienced significant toxicity changes after the ban increased their toxicity. However, this behavior was more pronounced in some subreddits and less evident in others. Overall, we observed that the findings for RQ1 in [13], which utilized the median, are similar to those presented in this section when using the mean and a causal inference method such as DiD.